# The potential applications of muography to revealing sea shipwrecks

### v3

@@ 2.3 Muon imaging of the underwater objects
As described in the previous section, the detector is positioned on the seabed, oriented toward the shipwreck with a slight inclination of approximately 10 degrees from the vertical. A detailed detector model was implemented to simulate muon hits for flux measurements. Each of three detector modules records the position of a passing muon, providing three spatial points for each event. Muon tracks are reconstructed by fitting a straight line through these three hit positions, from which the incident direction of each muon is determined. The reconstructed track yields the zenith and azimuth angles $(\theta,\varphi)$, defining the muon’s trajectory through the target. Accumulating many such tracks over time allows the detector to measure the angular distribution of the muon flux, which can then be used to calculate transmission and infer the internal density structure of the shipwreck.
@@
The MT technique involves two measurement configurations: with target and free-water (no target). The target configuration involves measuring the muon flux near or beneath the object of interest, where the water or material thickness between the detector and the surface constitutes the target. Angular fluxes are computed from reconstructed tracks, implicitly including acceptance and efficiency via the identical selection. The simulated transmission for a given column density $X$ along each line of sight is
$t(\varphi,\theta;X)=\frac{\Phi_{{target}}(\varphi,\theta;X)}{\Phi_{{notarget}}(\varphi,\theta)}$
@@
where no-target (free-water) configuration consists of ship filled with water extending to a depth of D meters, while the target configuration includes a ship filled with either crude oil or RDX. The result is a lookup table that maps column density $X$ to the corresponding simulated transmission $t(\varphi,\theta;X)$ for each angular bin. Finally, the measured transmission values are associated with their corresponding lines of sight and projected onto a 2D image plane, yielding a two-dimensional density map of the shipwreck. In practice, the detector is positioned on the seabed, oriented toward the target, and muon counts $N_{\mathrm{target}}(\varphi,\theta)$ are accumulated over the acquisition time $t_{\mathrm{acq}}$ for each directional bin. For each angular bin $(\varphi,\theta)$, invert the simulated model to determine the estimated column density:
$\hat{X}(\varphi,\theta)\quad\text{such that}\quad t(\varphi,\theta;\hat{X})\approx t(\varphi,\theta).$ (2.1)
@@
Subtracting the water baseline yields the excess or deficit in column density attributable to the shipwreck’s internal structure.
@@
For the projection onto a 2D image plane we treating the detector as a pinhole camera, the intersection point $\mathbf{p}_{kl}$ on the target plane for a muon arriving from direction $(\varphi_{k},\theta_{l})$ is:
$\mathbf{p}_{kl}=\mathbf{o}+R\,\hat{\mathbf{r}}(\varphi_{k},\theta_{l}),$ (2.3)
where $\mathbf{o}$ is the detector position. These points are projected onto the 2D image plane to assign $\Delta\hat{X}$ or $\bar{\rho}$ values to each pixel, creating a spatial density map. The spatial resolution depends on the detector’s angular resolution $\sigma_{det}$ and the multiple scattering angle $\theta_{0}$, which broadens the point spread function (PSF):
@@
where $x$ is the material thickness traversed, $X_{0}$ is the radiation length, and $p$ is the muon momentum.
@@ 3 Results
Figure 4(a) shows the angular distribution of muon counts simulated in the free-water configuration, with the shipwreck filled with water and the detector deployed at an underwater depth of 50 m. Figure 4(b) presents the difference in angular distribution of muons between two configurations where the shipwreck is filled with water in both cases. Figure 4(c) displays the one-dimensional projection of the difference histograms between these water-filled configurations. Figures 4(b) and (c) exhibit a uniform distribution, as both images correspond to the same configuration of the shipwreck filled with water, resulting in negligible differences. Figure 4(a) shows the angular distribution of muon counts simulated in the water-target configuration, with the shipwreck filled with crude oil. Figure 4(b) presents the difference (subtraction) of the angular distribution of muons between the water and crude oil configurations. Figure 4(c) displays the one-dimensional projection of the difference histograms between the water and crude oil configurations, highlighting the muon excess corresponding to the crude oil-filled region due to less absorption from its lower density. Figure 4(a) shows the angular distribution of muon counts simulated in the water-target configuration, with the shipwreck filled with RDX. Figure 4(b) presents the difference (subtraction) of the angular distribution of muons between the water and RDX configurations. Figure 4(c) displays the one-dimensional projection of the difference histograms between the water and RDX configurations, highlighting the muon deficit corresponding to the RDX-filled region due to greater absorption from its higher density.
Figure 2: Muon count distributions and differences for water - water-filled shipwreck configurations.
Figure 3: Muon count distributions and differences for water and crude oil-filled shipwreck configurations.
Figure 4: Muon count distributions and differences for water and RDX-filled shipwreck configurations.
The observed excess in muon flux in the x-projection of difference histograms (Figure 4), particularly at tangent values between -0.5 and 0.5, suggests a localized change in material density, here, a shipwreck filled with crude oil, which has a lower density than water.
The significant increase in muon flux observed in the x-projection of the difference histograms (water - crude oil configurations, Figure 4(c)) at $\tan(\theta)$ values of -0.5 $\div$ 0.5 can be attributed to the lower density of crude oil ($\approx 0.92\,\text{g/cm}^{3}$) compared to sea water ($\approx 1.03\,\text{g/cm}^{3}$), resulting in reduced muon attenuation and higher transmitted flux.
The significant deficit in muon flux observed in the x-projection of the difference histograms (water - RDX configurations, Figure 4(c)) at $\tan(\theta)$ values of -0.5 $\div$ 0.5 can be attributed to the lower density of RDX ($\approx 1.82\,\text{g/cm}^{3}$) compared to sea water ), resulting in reduced muon flux.
With sufficient angular coverage and longer exposure times, such 2D projections can be combined from multiple viewing positions to reconstruct a full three-dimensional tomographic model. For 3D reconstruction of shipwreck measurements in 4-8 positions around shipwreck are required.
@@ 4 Discussion
While muon imaging is still in its infancy for underwater applications, pilot studies show considerable promise. For instance, researchers at the University of Tokyo have successfully used muon tomography to image underwater volcanic structures, demonstrating its feasibility in marine environments. In the Baltic Sea, where wrecks such as the Som submarine or the so-called Baltic Sea Anomaly have attracted attention, muon imaging could help clarify the nature of obscured objects.
A particularly exciting application of underwater muon tomography is the study of ancient shipwrecks in the search for archaeological artifacts, such as gold or silver coins. Another potential motivation is the recovery of ancient lead, which has unique value as radiation shielding in low-background physics experiments of dark matter search due to its extremely low content of radioactive 210Pb. For these reasons, more detailed studies are needed to evaluate the performance of muon imaging in marine archaeology, with the goal of producing high-resolution 3D reconstructions of the internal structure of shipwrecks.
Advancements in compact, submersible muon detectors and AI-driven data analysis could make muon imaging more accessible for underwater archaeology. Integrating these technologies with existing systems, such as SeaTerra’s ScanFish or multibeam radar, could enhance wreck mapping efficiency. Collaborations between physicists, marine archaeologists, and exploration teams like Ocean X could further accelerate the adoption of muon imaging for investigating wrecks in the Baltic Sea and beyond, including the Mediterranean, North Sea, and Pacific Ocean, where similar preservation conditions exist.
@@ 5 Conclusion
In this study, we assess the feasibility of muon tomography as a non-invasive technique for detecting and characterizing submerged structures and hazardous materials within shipwrecks. Using Monte Carlo simulations, we modeled a simplified scenario of muon imaging of a trapezoidal shipwreck approximately 30 meters long, 15 meters high, and 15 meters wide, located at a depth of 50 meters. We considered three filling scenarios: the wreck fully filled with water, crude oil, or explosive RDX.
Assuming an exposure time of one week for a wreck located at a depth of 50 meters, our simulations demonstrate that muon imaging can resolve density contrasts sufficiently to distinguish between water, oil, and high-density materials, underscoring its feasibility as a practical tool for underwater hazard assessment.
To generate the underwater muon flux, we developed a muon generator that utilizes tabulated data simulated with GEANT4 and the CRY muon generator. Muons sampled over a 50 m × 50 m surface area using CRY were propagated through water to a depth of 35 meters with GEANT4, where they were detected by a sensitive detector that recorded event IDs, particle types (via PDG codes), kinetic energies, positions, and angular distributions. This tabulated data was then read by PrimaryGenerationAction.cc to generate muons originating from a 20 m × 20 m surface area underwater, facilitating the imaging of the shipwreck. The simulation results demonstrate that, with a measurement time of one week, oil and RDX can be definitively discriminated from water. Specifically, by subtracting histograms of angular distributions of detected muons for the water-oil scenario, an excess of muons due to the lower density of oil is observed. Conversely, a deficit of muons is detected in the case of the shipwreck filled with high-density RDX. These preliminary simulation results demonstrate the feasibility of applying muon tomography as a non-invasive tool for identifying and characterizing the internal contents of shipwrecks, including potentially hazardous materials, in challenging underwater environments such as the Baltic Sea.
Underwater muon radiography adapts muography principles to marine environments by accounting for seawater attenuation and employing detectors specifically designed for underwater operation. The integration of GEANT4-based underwater muon flux simulations with detector modeling enables realistic performance assessments, including expected track reconstruction rates for multi-module muon telescopes. By measuring the angular dependence of muon transmission, correcting for water depth, and comparing results with detailed simulations, it becomes possible to reconstruct 2D density maps of submerged structures. To reconstruct the ship’s structure and material distribution, the use of multiple muon detector locations on both sides is necessary. This configuration enables the creation of a 3D tomographic image of the interior structure, similar to what has been achieved in previous applications of muon imaging to old nuclear reactor [28]. This approach opens new opportunities for advancements in marine archaeology, geology, and underwater cultural heritage research.

### v5

@@ 2.2 Muon imaging of the underwater objects
In this study, we simulated a shipwreck as a trapezoidal object approximately 30 meters long, 15 meters high, and 15 meters wide, submerged at a depth of 50 meters to mimic the marine environment (as shown in Figure 5). The wreck was filled with crude oil (density = 0.92 g/cm3) or explosive hexogen (RDX) (density = 1.82 g/cm3), depending on the scenario.
To analyze the contents of shipwrecks, such as determining whether they are filled with water, hazardous crude oil, or explosives, the transport of muon flux through the target can be simulated. Underwater muons were generated using the PrimaryGeneratorAction class, which utilized tabulated muon data at a depth of 35 m and sampled over a 20 $\times$ 20 m2 surface. To achieve sufficient statistical accuracy, a total of 8 $\cdot~10^{9}$ muons were generated, corresponding to a scanning time of approximately seven days.
Figure 5: Schematic of muography imaging of an underwater shipwreck, modeled as a trapezoidal structure, with a muon detector placed on the seabed and oriented toward the shipwreck.
@@
By examining the attenuation of muon flux as it passes through different materials, the muon transmission (MT) technique enables the creation of two- and three-dimensional average transmission and density maps of the studied object, providing valuable insights into internal relative densities. Achieving a well-defined muographic image of large targets, such as underwater shipwrecks, often requires data acquisition periods extending to weeks or even months. This extended duration is necessary to accumulate sufficient muon statistics across various viewing angles, ensuring adequate resolution and contrast in the resulting density maps. The MT technique involves two measurement configurations: with target and free-water (no target). The target configuration involves measuring the muon flux near or beneath the object of interest, where the water or material thickness between the detector and the surface constitutes the target. Angular fluxes are computed from reconstructed muon tracks, implicitly including acceptance and efficiency via the identical selection. The simulated transmission for a given column density $X$ along each line of sight is
$t(\varphi,\theta;X)=\frac{\Phi_{{target}}(\varphi,\theta;X)}{\Phi_{{notarget}}(\varphi,\theta)}$ (2.2)
@@
where no-target (free-water) configuration consists of ship filled with water extending to a depth of D meters, while the target configuration includes a ship filled with either crude oil or RDX. The result is a lookup table that maps column density $X$ (in g/cm2) to the corresponding simulated transmission $t(\varphi,\theta;X)$ for each angular bin. Finally, the measured transmission values are associated with their corresponding lines of sight and projected onto a 2D image plane, yielding a two-dimensional density map of the shipwreck. In practice, the detector is positioned on the seabed, oriented toward the target, and muon counts $N_{\mathrm{target}}(\varphi,\theta)$ are accumulated over the acquisition time $t_{\mathrm{acq}}$ for each directional bin.
@@
In the inversion step, the measured transmission $t_{\mathrm{meas}}(\varphi,\theta)$ is compared with the simulated transmission $t_{\mathrm{simu}}(\varphi,\theta;X)$ to determine the reconstructed (inverted) column density $\hat{X}(\varphi,\theta)$. Subtracting the water baseline yields the excess or deficit in column density attributable to the shipwreck’s internal structure. For each angular bin $(\varphi,\theta)$, solve
$t_{\text{simu}}(\varphi,\theta;\hat{X})\approx t_{\text{meas}}(\varphi,\theta)$
for $\hat{X}(\varphi,\theta)$. The uncertainty is
$\delta X(\varphi,\theta)\approx\frac{\delta t_{\mathrm{meas}}}{\left|\partial t_{\text{simu}}/\partial X\right|}.$
Subtract the water baseline:
$\Delta\hat{X}(\varphi,\theta)=\hat{X}(\varphi,\theta)-X_{\text{water}}(\varphi,\theta).$
@@
The detector field of view is discretized into angular bins, indexed by $k$ and $l$, corresponding respectively to the azimuthal angle $\varphi_{k}$ and zenith angle $\theta_{l}$. For each angular bin $(\varphi_{k},\theta_{l})$, the reconstructed muon direction $\hat{\mathbf{r}}(\varphi_{k},\theta_{l})$ defines a ray originating from the detector center $\mathbf{o}$. Its intersection with the target (projection) plane, located at distance $R$, is given by
$\mathbf{p}_{kl}=\mathbf{o}+R\,\hat{\mathbf{r}}(\varphi_{k},\theta_{l}).$
The spatial resolution depends on the detector’s angular resolution $\sigma_{det}$ and the multiple scattering angle $\theta_{0}$, which broadens the point spread function (PSF):
@@
where $x$ is the material thickness traversed, $X_{0}$ is the radiation length, c is the speed of light, $\beta$ is the ratio of speed of the muon to speed of light and p is the momentum of the muon (in MeV/c).
@@ 3 Results
Figure 6(a) shows the angular distribution of muon counts simulated in the free-water configuration, with the shipwreck filled with water and the detector deployed at an underwater depth of 50 m. Figure 6(b) presents the difference in angular distribution of muons between two configurations where the shipwreck is filled with water in both cases. Figure 6(c) displays the one-dimensional projection of the difference histograms between these water-filled configurations. Figures 6(b) and (c) exhibit a uniform distribution, as both images correspond to the same configuration of the shipwreck filled with water, resulting in negligible differences. Figure 8(a) shows the angular distribution of muon counts simulated in the water-target configuration, with the shipwreck filled with crude oil. Figure 8(b) presents the difference (subtraction) of the angular distribution of muons between the water and crude oil configurations. Figure 8(c) displays the one-dimensional projection of the difference histograms between the water and crude oil configurations, highlighting the muon excess corresponding to the crude oil-filled region due to less absorption from its lower density. Figure 8(a) shows the angular distribution of muon counts simulated in the water-target configuration, with the shipwreck filled with RDX. Figure 8(b) presents the difference (subtraction) of the angular distribution of muons between the water and RDX configurations. Figure 8(c) displays the one-dimensional projection of the difference histograms between the water and RDX configurations, highlighting the muon deficit corresponding to the RDX-filled region due to greater absorption from its higher density. The significant deficit in muon flux observed in the x-projection of the difference histograms (water - RDX configurations, Figure 8(c)) at $\tan(\theta)$ values of -0.5 – 0.5 can be attributed to the higher density of RDX compared to sea water, resulting in reduced muon flux.
Figure 6: (a) Angular distribution of muon counts simulated for a shipwreck completely filled with water, with the detector at a depth of 50 m. (b) Difference in angular distributions of muons between two configurations, both with the shipwreck filled with water. (c) Projection of the 2D difference histogram in panel (b) onto the x-axis. The numerical values of the color maps are the same as those on the projection histogram in panel.
Figure 7: (a) Angular distribution of muon counts simulated for a shipwreck filled with crude oil, with the detector at a depth of 50 m. (b) Difference in angular distributions of muons between the water-filled and crude oil–filled configurations. (c) Projection of the 2D difference histogram in panel (b) onto the x-axis.
Figure 8: (a) Angular distribution of muon counts simulated for a shipwreck filled with water. (b) Difference in angular distributions of muons between the water-filled and explosive-filled configurations. (c) Projection of the 2D difference histogram in panel (b) onto the x-axis.
@@ 4 Discussion
With sufficient angular coverage and extended exposure times, combining 2D projections from multiple viewing positions enables the reconstruction of comprehensive three-dimensional tomographic models. This multi-angle approach facilitates systematic analysis of muon attenuation, scattering, and transmission properties, allowing for the characterization of the wreck’s internal structure with improved accuracy. When paired with detectors featuring high spatial and angular resolution, this configuration can produce detailed 3D images of internal features and estimate material densities with greater precision. Similar methodologies have been successfully employed in terrestrial applications; for example, muon imaging has been used to investigate the internal structure of decommissioned nuclear reactors, demonstrating the technique’s capability for high-resolution 3D characterization [33].
For underwater imaging, the detector’s spatial resolution is a critical factor influencing the accuracy of 3D reconstruction. In this study, a muon telescope was constructed using triangular extrusion scintillator bars, providing a spatial resolution of 2–3 mm as a cost-effective solution. However, if higher resolution is required for detailed reconstruction of the shipwreck’s internal structure, more advanced muon tracking systems with resolutions of 0.5–1 mm, such as those discussed in [34, 35], can be employed.
It is important to note that this study did not account for seasonal variations in muon flux or changes in water column density, as the primary objective was to evaluate the fundamental feasibility of the method. Incorporating such environmental factors in future studies could improve the accuracy of density estimates and enhance the robustness of the technique. Although these simplifications may introduce minor uncertainties in the reconstructed density values, they are unlikely to affect the overall assessment of the method’s potential. Muon generation in this work was performed using the CRY event generator, which tends to produce fewer muons (approximately 7–8×103 muons per m2 per minute) than the commonly referenced value of around 10,000 muons per m2 per minute. Consequently, the estimates presented here should be regarded as conservative. Additionally, CRY allows for adjustments based on geographic location, including the geomagnetic cutoff and solar cycle effects, ensuring that simulations can be tailored to specific sites.
Additionally, the thickness of the shipwreck hull was not simulated due to simplifications and computational limitations. This omission is justified, as 2 cm of steel corresponds to approximately 15.6 cm of water in terms of muon energy loss. Compared with the $\approx$15 m thickness of water or oil considered in the model, this introduces only a small systematic uncertainty (1-2%) in the muon flux, primarily affecting lower-energy muons, without significantly altering the observed differences between the water- and oil-filled configurations.
@@ 5 Conclusion
This study evaluated the feasibility of muon radiography as a non-invasive technique for detecting and characterizing submerged structures and hazardous materials within shipwrecks. Using Monte Carlo simulations, we modeled a simplified scenario of muon imaging for a trapezoidal shipwreck approximately 30 m long, 15 m high, and 15 m wide, located at a depth of 50 m. Three internal filling conditions were considered: water, crude oil, and the high-density explosive RDX. Assuming a one-week exposure time, the simulation results indicate that muon imaging can resolve density contrasts sufficient to distinguish between these materials, demonstrating the method’s potential as a practical tool for underwater hazard assessment.
These preliminary findings confirm the feasibility and promising potential of muon radiography for non-invasive identification and characterization of shipwreck contents in challenging underwater environments such as the Baltic Sea, while acknowledging that further refinements, such as accounting for environmental variability and optimizing detector performance, are essential for advancing practical applications.
Beyond environmental and safety applications, this approach offers promising opportunities in marine archaeology, geology, and underwater cultural heritage research. By enabling detailed, non-destructive imaging of submerged structures and geological formations, muon radiography can enhance our understanding of underwater sites, support preservation efforts, and reveal historical and geological insights that are otherwise difficult to obtain.